A 10 g bullet having a charge of 4 micro coulomb is fired at a speed of 270 m/s in a horizontal direction. A vertical magnetic field of 500 micro tesla exists in the space. Find the deflection of the bullet due to the magnetic field as it traveld through 1 m. Make appropriate approximations.

A 10 g  bullet having a charge of 4 micro coulomb is fired at a speed of 270 m/s in a horizontal direction. A vertical magnetic field of 500 micro tesla exists in the space. Find the deflection of the bullet due to the magnetic field as it traveld through 1 m. Make appropriate approximations.

Grade:11

1 Answers

Arun
4 years ago
Dear Abhishek

F=qvB=ma   ;    a=qvB/m=54micro units(mtrs./sec.2)

S=1/2at2  where t=100/270     ;  S=100/27 micro mtrs.  :this is the deflection.